FEL Upgrade Project Weekly Brief
April 15-19, 2002


We are pleased to acknowledge receipt of a valuable piece of equipment that will benefit the UV FEL Upgrade program. Brookhaven National Lab has shipped to us a prototype version of the Undulator A that was built by a Cornell University-Argonne National Lab collaboration. The undulator was most recently carefully characterized during use in the Brookhaven" High Gain High Harmonic Experiment". This equipment is being loaned to the JLab FEL Program on a long term loan arrangement. We thank our colleagues at Cornell and Brookhaven for the valuable loan and useful characterization data.

Lots of construction progress this week: among the highlights the first girder assemblies for the e-beam transport system will be assembled this weekend.

FEL management attention this week was devoted to a careful analysis of the remaining budget for completion of the 10 kW upgrade project and the commissioning activities. The team continued to prepare and review presentation material for the May 1-3 semiannual review.

A revised "Work for Others" contract for the FY02 increment in funding for the UV FEL project was forwarded to our AFRL contract monitor from the DOE JLab Site Office. This contract revision would add an additional $1.5M to the work scope for the JLab/Aerospace collaboration on the UV FEL project.

An article was published this week in the May/June issue of the "Industrial Physicist" that summarizes materials science applications of the FEL and the upcoming "Industrial Physics Forum" which will be hosted by JLab in October 2002. The article can be found on the Web at:


Gwyn Williams attended a DOE sponsored workshop this week in Charlotte on the "Genomes to Life" program which included discussions of potential applications of the planned synchrotron addition to the FEL Facility.

On a less positive note, we found out this week that the anticipated funding for the $3M addition to FEL facility to house the Helios synchrotron did not make the FY03/04 Virginia state budget approved on April 17th by the General Assembly. We have been told by the Governor's Office that the administration intends to restore the budget item during the next General Assembly session in January 2003.

WBS 3 (Beam Physics):

Dipole field characterization review suggests dipoles (with shaving) can provide operationally acceptable core field & central field integrals at energies we'll be able to reach for the next while. AES is characterizing behavior with shims - we'll see what 210 MeV/c looks like soon.

Moving on to field homogeniety, we've projected the beam stay-clear onto AES results - figures modified from Tom Schultheiss's results follow - these show the field variation across the beam is pretty much at the 10^-4 level we need.

Figure 1: Beam stay clear (dashed red) and central orbit (solid red) superposed on AES (Schultheiss) contour
map of GX field

Figure 2: Beam stay clear projected onto AES results for field (T) vs. transverse position of straight-line paths through GX magnet at 80 MeV/c

Cogitation on commissioning continues - perusal of IR Demo log entries from commissioning suggests a six month effort is reasonable provided an adequate number of victimizable participants are available to run 2 shifts/day, 5 days/week (order 10 experienced accelerator physicist/engineering staff).

WBS 4 (Injector):
Supported W&M implant efforts and the first test tube run was started to see if arcing problems had been resolved. Continued to work on electrode polishing. The BPM/ bellows after the lightbox was resolved. The BPM will be manufactured and tested.

Gun HVPS - The drawings for the Gun HVPS Tank as a pressure vessel were provided by the contractor and reviewed by the JLab engineers. The tank drawings were approved with minor corrections. The construction can now proceed with a May delivery planned.

WBS 6 (RF):
Quarter HVPS - Work continued on unit 4. The control wiring is about 60% complete. More meters were received and installed this week. A different PLC is being ordered to simplify the communication task with EPICS. A major portion of this communication software is already written. Next week the welders should be able to return to the new LCW piping effort. Thermal switches were removed from the SCR Controller and returned to the vendor for replacement. We needed the N.C. type contacts for our machine safety interlocks.

Quarter Klystrons - The last of the plumbing parts are due early next week. This will allow the klystron carts to be finished by the Machine Shop.

WBS 8 (Instrumentation):
Cable installation began in earnest this week, all 45 beam viewer signal cables have been pulled into the vault to their respective destinations. The remaining (ripout) beam viewers; ITV0F02, IMS0F03, ITV0F04, IXM0F06, IMS01F01 were modified in place with the new interface box and tested successfully. Thanks to Jim and his army of students! In one of Mark Augustine's first official acts, post re-org, he and his crew muscled our new trim rack into the FEL building - great job! Also the dipole string 'box' power supply was moved into its location. The field wiring will begin next week. The installation activity will be 'taking off' since the >month long accelerator shutdown is complete.

The beam viewer control chassis were installed into the racks at FL03B09, the debugging of the software screens and field termination will begin next week. The system print for the beam viewer system has also been started. This is one of the more strait forward tasks to be done so we're knocking it off first.

The Machine Protection System (MPS) and Drive Laser Pulse Controller (DLPC) design discussions made good progress this week. With the new VME based Beam Loss Monitor (BLM) designs complete Rich shifted gears to the drive laser & MPS controls. A cut was done at the different modes of machine operation (tune-up, IR, UV), and how the few hundred devices (viewers, vacuum valves, BLMs laser alignment mode, gun HV...) will all be integrated into safe and seamless operations. The Altera (gate array) timing boards (24) were received and are ready for assembly. The VME interface and MPS input daughter boards are in final assembly and will be tested starting next week.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
• At Magnet Enterprises International (MEI) in Oakland CA. The fifth coil looks very good. The 6th is in potting.
• Progress on the cores is slow. Return legs are complete. Contrary to last week’s report, the grinding of the batch of 4 top and 4 bottom slabs for the IR machine is not complete. Delivery slipped a week so far and may still not be ready. The promise by the follow on machining subcontractor that this batch will be machined "by the end of the month" is no longer valid. It will not be ready for the review. We are issuing a "notice of concern" to MEI about their late delivery using these sub contractors.

Injector Dipoles (DU/DV)
• Master Machine is nearly finished rough machining the slabs for the yokes. Heat treating will start next week. Field clamps are finished. Mu metal sheets for the GU did not come in to the uniform thickness specified and we are working with the supplier to correct that.
• WANG NMR is ready to pot all the coils.
Arc Dipoles (GY, GX, GQ)
• Process Equipment Co. received the steel for GQ-GX magnet cores. They are preparing their shop routing sheets, rough machining material and have about one more week of parts ordering to do
• Wang NMR is working on the winding fixture for the coils.
• AES has completed their first pass at magnetically modeling the GQ-GX dipoles at lower energy with shorter pole tips to match David Douglas’ use matrix. Their results are being studied.
• Wang NMR started winding the first GY coil. It is slow going with this difficult coil. They delayed in starting impregnation of the GG coils by other work.
• Bosma Machine has the two bottom GY core pieces done and is ready to paint. The top plates will be finished planning this week. Return legs are complete. Good progress.
3 inch quad (QX)
• Further testing was stopped by (1) failure of the power supply while doing "bang-bang" hysteresis curves and (2) bad bearings in the probe drive motor. At week’s end we have located three alternate supplies and bearings were being shipped. Eight more of the UV quads were received.
Trim Quad (QT)
• Milhous Control of Virginia reports that a first article magnet may be finished by next Wednesday.
Sextupole (SF)
• DULY Research is working on completing drawings of the magnet. We are down to discussing tolerance schemes to apply to the pole tip drawing, routing schemes for conductors between coils as well as cooling routing.
Octupole (OT)
• Work on this magnet is on furlough until the sextupole is designed.
Beam Line and Vacuum
• Design of the ARC Chambers is complete to the JLab checking stage.
• The assembly and detail drawings of the girders for the six quadrupole telescope in front of the first arc was signed off.
• The second and almost complete listing to stand positions is in checking before being given to the Survey and Alignment Group. They plan to mark floor positions when they return from their work in CEBAF during its down period.
• Arc dipole stand fabrication is nearly complete while the optical chicane dipole stands were signed and sent to procurement.
• Girder assembly will start over the weekend.

WBS 10 (Wiggler):
We took final measurements on the wiggler this week. The first measurements showed that the wigglers had the wrong polarity for X-ray work. The leads were reversed and the field checked. The corrector strengths were the same as before to within measurement precision so changing the polarity has no effect on the wiggler calibration. The interaction of the wiggler and dispersion section lead to a reduction in the strength of the last half pole in the wiggler as expected and a slightly reduction in the outer pole of the dispersion section (not expected but a small effect). The net result of this is that the wiggler corrector coils will have to run about 1.5% stronger than previous calibrations indicated. Otherwise no problems seen. Took some measurements of the noise offsets and started staging for vacuum chamber reinstallation. We are on schedule for moving the wiggler to the vault next Wednesday.

WBS 11 (Optics):
We received the bids for the optical cavity internals. The low bid was only 3% over our estimate. The OCMMS hardware fabrication procurement cost came in ~ 2x higher than expected, until we discovered the fabricator estimated the cost of only one unit, when in fact we needed two. That procurement was also awarded. The procurement for additional vacuum valves for the optical transport system was signed. The design effort for the optical cavity vacuum vessels is making good progress. Modeling of the 50kW laser beam dump is complete and prototypes are being made in-house. A model for the broadband optical cavity was developed and is being analyzed. Design of the chamber and optics for the THz light pickoff continue.

This week, we had a breakthrough on the drive laser pulse compressor. After using a beam profiler to confirm that the drive laser mode quality was still high, more time was spent with the mode-matching lenses that couple beam into and out of the fiber. A set was finally found that didn't produce diffraction rings. The next step is to finish aligning the fiber output into the rest of the optics and measure the pulselength of the compressed pulse. Work continues on the OCMMS piezo transducer controls. Raising the mechanical stiffness of the many (~ 6) transport mirror mounts raised the lowest mechanical resonance to a value above 100Hz. A 2nd order Bessel filter (soon to be a 4th order) is being evaluated to reject these spurious signals. The filter introduces phase shift, which must be characterized so we can introduce the appropriate correction at the inverter. We are testing the linear actuators for the optical cavity mirrors to confirm they produce rated thrust before stalling.

Other activities:
We did more preparatory work to prepare for the May 3rd review.

We are testing a technique for metallizing the sides of sapphire mirrors to facilitate the use of a heat-transfer medium. As we continue the design of the optics, we are tabulating values of thermal and optical properties of sapphire and calcium fluoride, our lead candidates for outcoupler substrates. In doing this, we're finding that there is a paucity of loss data on sapphire in the ultraviolet. We probably will need to make these measurements to better characterize the material.